If a transmitter is connected to a cable that has a loss of 6 db and each connector (one on the transmitter end of the cable and one on the antenna end of the cable) has a loss of 1 db, what would be the total loss in db?
Answer #1

The cable which is disscused in question is coaxial cable.

Here, Number of connectors of cable = 2

one of the cable end has loss = 1db

Therefore, signal loss due to both connectors = 2 * 1 = 2db

Signal loss due to cable = 6db

Hence Total loss due to cable and connectors = 2 + 6 = 8 db